### v1

# Layered Insights: Generalizable Analysis of Authorial Style by Leveraging All Transformer Layers

Abstract
We propose a new approach for the authorship attribution task that leverages the various linguistic representations learned at different layers of pre-trained transformer-based models. We evaluate our approach on three datasets, comparing it to a state-of-the-art baseline in in-domain and out-of-domain scenarios. We found that utilizing various transformer layers improves the robustness of authorship attribution models when tested on out-of-domain data, resulting in new state-of-the-art results. Our analysis gives further insights into how our model’s different layers get specialized in representing certain stylistic features that benefit the model when tested out of the domain.
**footnotetext: These authors contributed equally to this work
1 Introduction
Authorship attribution (AA) is the task of analyzing the writing style of pairs of texts in order to predict whether they are written by the same author. Approaches developed for this task either focus on explicitly modeling linguistic attributes of texts Koppel and Schler (2004) or rely on pre-trained transformer-based models to learn relevant features to the task from big training corpora Rivera-Soto et al. (2021); Wegmann et al. (2022). The latter approach achieves state-of-the-art results on the AA task. However, it focuses on modeling texts by only learning from the final output layer of the transformer, ignoring representations learned at other layers. A large body of research has been investigating the inner workings of pre-trained transformer-based models Rogers et al. (2020), showing that different layers are specialized in modeling different linguistic phenomena of texts, with lower layers mainly modeling lexical features while higher ones model sentence structure Rogers et al. (2021).
Figure 1: Our approach is a transformer-based model where the initial embeddings (Init. Embed.) and the output of each layer passes through a projection layer, obtaining N+1 document representations that are jointly optimized on the authorship attribution task using contrastive learning
In this paper, we hypothesize that leveraging the representations learned at different layers of a pre-trained transformer can lead to more effectiveness on the authorship attribution task, especially since the AA task requires modeling similarities between different linguistic aspects of the texts. For example, the writing style of an author can be realized by their use of similar vocabulary (lexical), sentence structure (syntax), or even the overall discourse they follow in their writing. To study this hypothesis, we develop a new approach to the task that learns to capture these different stylistic aspects by leveraging all the transformer layers. In particular, as highlighted in Figure 1, starting from a pre-trained transformer-based model with N layers, our approach projects the initial token embeddings as wel as each of the N layers’ embedding into a new embedding space, where we apply a contrastive loss that refines this space to better capture the similarity between texts with respect to the original representation. The model is then trained jointly by averaging the contrastive losses at all the N+1 projections. During prediction, given two texts, the analysis of whether they belong to the same author can be decomposed into comparing their similarities at these different learned projections, leveraging more robust predictive power by aggregating these scores.
To evaluate our hypothesis, we follow the same evaluation setup of Rivera-Soto et al. (2021). We train and evaluate our approach against state-of-the-art models on three datasets: Reddit comments, Amazon reviews, and Fanfiction short stories. Our experiments demonstrate that, for the out-of-domain scenario, our approach performs best in all settings, with the highest increase of around 50% in the mean reciprocal rank (MRR) for our model trained on the Fanfiction dataset and evaluated on Reddit. Results, however, in the in-domain scenario are mixed. These insights indicate strong generalization capabilities of our model compared to baselines, which we argue is more relevant in practical scenarios of the AA task than in-domain. This also provides evidence of the importance of leveraging the various layers of the base transformer when working on the AA task. Furthermore, we perform experiments to analyze the role each layer plays in the overall performance of our model. We found that certain layers became more relevant when our model tested in the out-of-domain setting, leading to high performance.
In summary, our contributions are the following:
  • A state-of-the-art approach for authorship attribution in out-of-domain scenarios.
  • Empirical evidence supporting the generalizability of our approach compared to baselines
  • Insights into the different representations learned in our model
2 Related Work
Transformer-based models have shown strong advances in learning numerous natural language processing tasks. In the pre-training phase, these models are trained on large corpora to predict the next token given a context window, leading them to acquire strong text representations that can be leveraged to learn downstream NLP tasks. In a comprehensive review, Rogers et al. (2021) showcases how the internal representations of transformers become increasingly specialized across layers, with lower layers capturing surface-level syntactic information and higher layers encoding deeper semantic and task-specific knowledge. van Aken et al. (2019) highlights how specific layers in transformers contribute differently to question-answering tasks, underscoring the model’s specialization at varying depths. Additionally, Fan et al. (2024) provided evidence that only a subset of layers is required for specific downstream tasks, emphasizing the potential for task-dependent optimization of computational resources. These studies collectively emphasize that linguistic knowledge is distributed across layers in a hierarchical and task-dependent manner. Inspired by this literature, in this paper, we study the task of authorship attribution at these different levels of granularity.
Author Attribution. Traditional methods rely on manually designed features, such as syntactic patterns and function word frequencies, to capture distinctive writing styles Koppel and Schler (2004). With the rise of deep learning, transformer-based models have set new records by leveraging their ability to capture nuanced stylistic patterns in text. For instance, Rivera-Soto et al. (2021) employed a contrastive learning approach to map texts written by the same author into similar regions of an embedding space. Similarly, Wegmann et al. (2022) introduced methods to ensure that embeddings focus on style rather than content, allowing for more reliable attribution. Despite their success, these approaches rely only on the final layer of the transformer model as the source of text representation. In contrast, our approach leverages different text representations from all the transformer’s layers.
3 Approach
Given mounting evidence from the literature on the various linguistic capabilities of the different layers of transformer-based models, we propose a multi-layer contrastive learning approach that explicitly models authorship attribution at varying levels of linguistic granularity. Instead of relying on a single representation derived from the final layer, our approach extracts embeddings from all the N layers of the transformer to leverage its linguistic representations for the authorship attribution task.
As highlighted in Figure 1, our model is RoBERTa-based Liu (2019) with additional N+1 Feed Forward layers that project each of the hidden states of RoBERTa into a new embedding space, including the initial token embeddings. We then apply a contrastive loss function on each of these N+1 embedding spaces to pull documents written by the same author together while keeping other documents further apart in the corresponding space. At each training step, the average loss of all the N contrastive losses is computed and back-propagated throughout the model. We hypothesize that this architecture ensures that stylistic features across multiple linguistic levels contribute to the final representation, allowing the model to learn distinct and complementary signals from different layers. Rather than collapsing stylistic information into a single embedding, we maintain separate representations for each layer. In Section 5.1, we present an analysis confirming how the various learned layers of our trained model capture different linguistic representations from texts. During inference, given two texts, we compute their similarity at each learned layer projection. These layer-wise similarity scores are aggregated to produce a final similarity measure that determines the likelihood of these two texts being written by the same author.
4 Evaluation
We first describe the experiment setup to evaluate our model’s performance on authorship attribution, then present our results.
Training Dataset
Evaluation Dataset Model Reddit Amazon Fanfic
R@8 MRR R@8 MRR R@8 MRR
Reddit MultiLUAR 67.87 53.31 26.92 18.03 21.11 14.32
LUAR 69.67 54.78 22.67 14.65 14.61 9.49
Conv 56.32 42.38 6.30 9.70 5.74 3.90
Tf-Idf 10.34 6.77 7.65 5.03 6.97 4.63
Amazon MultiLUAR 72.63 60.74 84.68 72.65 56.69 44.36
LUAR 71.84 58.51 82.54 68.99 37.12 26.39
Conv 60.20 47.60 74.30 60.06 34.90 25.90
Tf-Idf 43.70 35.50 31.61 24.86 21.46 16.45
Fanfic MultiLUAR 43.26 33.96 39.21 29.98 51.78 44.33
LUAR 42.50 32.30 34.56 25.04 55.38 45.07
Conv 40.66 30.98 24.99 17.98 47.98 39.02
Tf-Idf 25.22 18.72 26.04 19.37 31.37 22.53
Table 1: Evaluation results comparing our approach (MultiLUAR) against other baselines on the three datasets for in-domain (highlighted grey) and out-of-domain settings. The highest values are marked bold.
4.1 Experiment setup
In our evaluation, we closely follow the experiment setup of Rivera-Soto et al. (2021) to reproduce their results and compare our approach against theirs.
Task.
For the authorship attribution task, the input is two collections of texts, each written by a single author, and the output is a score reflecting the likelihood of these two collections being written by the same author.
Datasets.
We evaluate the effectiveness of our model on the Reddit, Amazon and Fanfiction datasets used in Rivera-Soto et al. (2021). These datasets represent distinct domains with different linguistic styles, levels of formality, and authorial consistency. Each dataset is partitioned into training and evaluation, ensuring that evaluation texts are temporally disjoint from training data where timestamps are available Andrews and Bishop (2019). For Amazon, we use 135K authors with at least 100 reviews, training on 100K, and evaluating on 35K Ni et al. (2019). The Fanfiction dataset Bevendorff et al. (2020) consists of 278,169 stories from 41K authors, with evaluation restricted to 16,456 authors, each contributing exactly two stories Bischoff et al. (2020). The Reddit dataset includes 1 million authors, with training data from Khan et al. (2021) and evaluation using a disjoint, temporally future set Baumgartner et al. (2020). Further information regarding the datasets can be found in the appendix.
Metrics.
We evaluate all models using Recall-at-8 (R@8) and Mean Reciprocal Rank (MRR). Recall-at-8 measures the probability that the correct author appears among the top 8 ranked candidates, while MRR evaluates ranking quality based on the position of the first correct author match. Higher values in both metrics indicate stronger performance Voorhees and Tice (2000).
Baselines.
We compare our approach against the following baseline models trained on the same datasets. The primary baseline is the model proposed by Rivera-Soto et al. (2021), called LUAR. This baseline is a contrastive learning-based authorship attribution model with a RoBERTa model (paraphrase-distilroberta checkpoint [^1] https://huggingface.co/distilbert/distilroberta-base) as its base transformer. Given the two text collections, the model samples excerpts of 32 tokens and passes them through a RoBERTa transformer, obtaining only the final layer representation of each excerpt. An attention mechanism and max pooling layer are applied to aggregate the extracted representations into a single embedding representing each text collection. A contrastive loss is then applied to these representations to make them closer for positive pairs (written by the same author) and further apart for negative pairs. Additionally, we evaluate against the same baselines presented in Rivera-Soto et al. (2021). The first is a convolutional model Andrews and Bishop (2019), which encodes text representations using subword convolutional networks rather than transformers. The second baseline is a TF-IDF baseline, which represents documents as bag-of-words vectors weighted by term frequency in the text and normalized by inverse document frequency. We follow the training process of Rivera-Soto et al. (2021) to train a reproduced version of their LUAR model while we report the performance of the other two baselines as mentioned by the authors.
Our Model.
Our model extends the original LUAR architecture mentioned above by utilizing multiple transformer layers instead of a single final-layer embedding. Each layer’s hidden states are projected into separate embedding spaces and optimized using contrastive loss, capturing stylistic features across different linguistic levels. As a result, our model has seven projection layers: Layer 0 is the projection of the initial token embeddings of paraphrase-distilroberta transformer [^2] We also experimented with excluding the initial token embeddings, but the performance is worse. See Table 3 in the appendix, and layers 1 to 6 are the projection of its six hidden layers. We refer to our model as MultiLUAR. Following the LUAR training paradigm, we train a model for each of the three datasets for 20 epochs, respectively, with validation after every five epochs. We use a batch size of 84 and sample from every author 16 text excerpts of 32 tokens each. We evaluate two scenarios: in-domain (same dataset for training and evaluation) and out-of-domain (trained on one dataset, evaluated on other out-of-domain datasets). This setup allows us to assess the model’s generalization ability across different writing styles and domains, a scenario that has proven to be harder to achieve strong results in. Further details on the training process can be found in the appendix.
Training Dataset Evaluation Dataset Model Reddit $\cup$ Fanfic Reddit $\cup$ Amazon Amazon $\cup$ Fanfic R@8 MRR R@8 MRR R@8 MRR Reddit Ours 64.13 49.46 63.96 49.44 29.95 20.62 LUAR 56.35 41.20 60.58 45.46 20.40 13.11 Amazon Ours 69.43 57.04 89.00 78.38 85.99 74.23 LUAR 53.51 39.63 84.84 72.09 79.60 64.83 Fanfic Ours 54.16 44.56 43.24 33.54 51.87 42.63 LUAR 57.51 46.32 40.69 30.49 51.84 41.79
Table 2: Evaluation Results after training on multiple domain datasets
4.2 Results
Similar to Rivera-Soto et al. (2021), we examine the performance of our model when trained on single as well as multiple domains. We present our findings in the following sections.
4.2.1 Single-Domain Training
We first evaluate our MultiLUAR model in the single-domain setting, where models are trained on a single dataset and tested on the test split of all three datasets to assess their generalizability. Table 1 shows the recall at 8 (R@8) and mean reciprocal rank (MRR) scores of all models trained and evaluated on the different datasets. In general, LUAR and our MultiLUAR model outperform the other two baselines in all scenarios. However, for in-domain evaluation, we observe that our approach performs better than LUAR only on the Amazon dataset, while it performs slightly worse in the case of Reddit and Fanfiction datasets. This suggests that for in-domain evaluation, final-layer representations contain sufficient information for authorship attribution, and incorporating multiple layers might lead to a slight degradation in the performance. We argue that the improvement on the Amazon dataset is due to the structured writing style allowing the different layers to capture more nuanced linguistic representations from texts. In contrast, informal and discussion-based datasets, such as Reddit, exhibit less improvement, likely due to the topic-driven nature of the writing, which can overshadow stylistic signals.
For out-of-domain evaluation, our MultiLUAR approach demonstrates substantial improvements over LUAR baseline, significantly enhancing cross-domain generalization and establishing new state-of-the-art results. MultiLUAR trained on homogeneous datasets, such as Amazon, show remarkable adaptability to free-form writing (An MRR of 18.03 for MultiLUAR compared to 14.65 for LUAR on Reddit dataset), while those trained on Reddit generalize exceptionally well to both Amazon and Fanfiction. Compared to LUAR, our model substantially improves transferability, highlighting the importance of leveraging different linguistic granularity levels rather than relying solely on final-layer representations. Most notably, in the Fanfiction transfer setting, where LUAR struggles, our approach achieves 50% R@8 improvements, effectively capturing deep stylistic signals that enhance authorship attribution even in highly creative writing domains. These results confirm that our approach enhances robustness, making our approach more effective for authorship attribution across diverse linguistic contexts.
4.2.2 Multi-domain Training
Additionally, we replicate the same experiment in Rivera-Soto et al. (2021) where we train LUAR and our MultiLUAR model on two domains and evaluate them on the third domain to examine whether exposure to multiple domains enhances MultiLUAR’s robustness vs LUAR. To this end, we train our approach on dataset combinations and evaluate them on individual domains. We train models on domain pairs using mini-batches of 256 that are randomly selected from document collections, ensuring equal representation from both domains. Since authors are disjoint across domains, shared authorship occurs only within the same domain.
Figure 2: The linguistic capability of each of the 6 layers of paraphrase-distilroberta model (right side) and MultiLUAR trained on Reddit dataset (left side), as probed using Holmes explorer.
Table 2 presents the results. Overall, our MultiLUAR model benefits more than LUAR from including different training domains, achieving higher scores in almost all cases. Comparing with the single-domain results shown in Table 1, we find that introducing Reddit data consistently improves transferability, reinforcing its diverse linguistic coverage as an essential factor in cross-domain generalization. Unlike LUAR, where adding Fanfiction dampened the performance, our MultiLUAR model enhances out-of-domain generalization because it effectively leverages different linguistic granularity levels, allowing it to capture both fine-grained lexical choices and higher-level stylistic structures. This enables the model to distinguish authorial identity beyond surface-level patterns, improving performance across structurally diverse domains. This is particularly evident in models trained on Amazon $\cup$ Fanfiction; where LUAR failed, MultiLUAR has a strong performance on Reddit by integrating multiple stylistic cues. Conversely, introducing Amazon leads to mixed results, improving transferability in some cases while slightly reducing in-domain performance. Furthermore, introducing Reddit into in-domain evaluations also leads to performance improvements. Given its strong transferability, we observe enhanced in-domain performance across other datasets as well. Notably, due to MultiLUAR’s significant improvements in Fanfiction’s out-of-domain capabilities, we also see in-domain gains when paired with Fanfiction, particularly in the Amazon $\cup$ Fanfiction setting.
5 Analysis
In this section, we take an analytical look at what our model’s projection layers capture linguistically and what role they play in the final prediction in both in and out-of-domain scenarios.
[…]
5.1 Learned Linguistic Representations
As mentioned, research showed that during the pre-training phase of transformer models, their layers acquire different linguistic capabilities. In this section, we look at this phenomenon in our trained MultiLUAR model to inspect whether the training process on authorship attribution task leads to different layers capturing different linguistic phenomena. In particular, we use the Holmes benchmark Waldis et al. (2024), which consists of over 200 datasets and different probing methods to evaluate the linguistic competence in transformers, including syntax, morphology, semantics, and discourse at different layers. We run the probing method of this library over paraphrase-distilroberta, the base transformer of our model, quantifying the linguistic competence of its initial embedding (layer 0) and all its 6 layers. We then compare it to MultiLUAR model that we trained on the Reddit dataset.
Figure 4: Percentage of text pairs (y-axis) having significantly high similarity score w.r.t. different layers (x-axis) of MultiLUAR trained on (a) Reddit, (b) Amazon, and (c) Fanfiction. These percentages are broken by the evaluation dataset: Reddit, Amazon, and Fanfiction
Figure 2 visualizes the linguistic competence of all layers for both models. As for RoBERTa model (Figure 2.b), we observe that for all four linguistic phenomena, layer 3 has the strongest representation rather than the final layer (a 0.5 score in Morphology and around 0.7 on Semantics, Discourse, and Syntax). While layer 6 is better at capturing semantics and morphology than layer 1 (0.6 and 0.38 compared to 0.5 and 0.3, respectively), it is worse at discourse and syntax. Since the authorship attribution task relies on comparing text similarity under different linguistic aspects rather than only semantics, these insights support the need to consider all transformer layers when solving the authorship attribution task. Furthermore, we can see that all layers are stronger at syntax, discourse, and semantics compared to morphology. Now, in comparison, looking at Figure 2.a that presents our MultiLUAR model’s layers linguistic capacity, we can observe the following. The competence of all layers is relatively stronger for morphology and syntax compared to discourse and semantics. This is an indication that syntax and morphology are more important for the task of authorship attribution than semantics and discourse, at least on the Reddit dataset that the model is trained on. Furthermore, we can see that different layers become specialized in specific linguistic phenomena. For example, layer 4 represents morphology best; layer 3 captures syntax, and layers 0 and 6 capture discourse better than other layers. One can also notice that, overall, MultiLUAR model’s linguistic capacity is lower than paraphrase-distilroberta across all layers. For example, at morphology, the highest score of MultiLUAR is 0.35 compared to 0.5 for RoBERTa. This decrease is expected since these projection layers of our model have a lower dimension size (512 compared to 768).
5.2 Layer to Prediction Contribution
We conduct experiments to analyze our model’s behavior through an ablation study and by quantifying each layer’s significance to the final prediction.
5.2.1 Ablation Study
The final prediction of our model is an aggregation score of the cosine similarity between all layer representations of the input pairs. Therefore, our ablation study evaluates the impact of different layers in the MultiLUAR model by systematically removing each layer from the aggregation and measuring its effect on attribution performance across three datasets: Reddit, Amazon, and Fanfic, allowing us to categorize this contribution for in and out-of-domain scenarios. Figure 3 presents the ablation results as a bar chart for each MultiLUAR model when evaluated on each of the three datasets while removing each of its layers from the predictions. Each bar reflects the change in the model’s performance in terms of R@8 after removing a specific layer. Negative bars indicate that the layer harmed the model’s performance.
MultiLUAR trained on Fanfic
Looking at the final column of Figure 3, we notice that when initial layers (Layers 0-3) are removed, there is a significant performance drop in the out-of-domain scenario (Figure 3.2.3 and 3.1.3). This suggests that the lower layers capture stylistic features that generalize well across different datasets. However, this comes with a cost, these initial layers contribute negatively when the model is evaluated in an in-domain setting (Figure 3.3.3), and removing them leads to an increase in the performance. In contrast, the latter layers (4, 5, and 6) contribute more to the in-domain setting; removing them harms the model’s performance (Figure 3.3.3).
MultiLUAR trained on Amazon
In the second column of Figure 3, we observe a relatively similar behavior. The final layers (4 to 6) contribute mostly to the in-domain setting (Figure 3.2.2). Layers 1 to 3, while benefiting the performance on Reddit (out-of-domain), contribute negatively to in-domain – removing them improves the performance on the Amazon dataset (Figure 3.2.2).
MultiLUAR trained on Reddit
Now, we consider the first column of Figure 3. We observe that removing layer 0 leads to performance degradation for Amazon and Fanfic (out-of-domain) while contributing negatively to the in-domain scenario. Layer 6 then has the highest contribution for the in-domain scenario (Figure 3.1.1).
Overall, the findings reinforce the importance of multi-layer representations in author attribution tasks. Unlike prior work that relies solely on final-layer embeddings, our results demonstrate that early and middle layers’ embeddings contribute significantly to out-of-domain generalization, with various layers playing different importance roles based on the evaluation dataset’s style characteristics. Higher layer embeddings then capture more nuanced, domain-specific stylistic cues. The MultiLUAR approach, which aggregates information across all linguistic granularity levels, provides a more robust framework for modeling author writing style. These results highlight the limitations of single-layer methods and emphasize the necessity of incorporating multiple levels of linguistic representation for improved attribution performance.
5.2.2 Layer significance
We analyze how relevant a specific layer is for predictions on each test dataset (Reddit, Amazon, and Fanfiction). We quantify layer relevance by computing the percentage of instances in the test set (a pair of author texts) with a cosine similarity higher or lower than the average score by a given margin. Specifically, given a corresponding dataset, a MultiLUAR model, and one of its layers, we sample a set of 5k text pairs, measure the cosine similarity of their representations at all layers, and compute the percentage of cases where the z-score of the corresponding layer is higher than 1.5 for positive pairs (indicating cosine similarity 1.5 standard deviations higher than average) or lower than -1.5 for negative pairs (indicating lower cosine similarity than average). We visualize these percentages in Figure 4 for the three MultiLUAR models trained on Reddit, Amazon, and Fanfiction datasets. Looking at Figure 4, we can see an overall pattern where the final layer plays a more specific role in predicting in-domain cases while early layers of the model contribute to the predictions of the out-of-domain cases, which aligns with our ablation study presented earlier. For example, considering at MultiLUAR trained on Reddit (Figure 4.a), we observe that layers 0 and 6 are mainly relevant to the Reddit instances, while the other layers, such as layer 1, play a more prominent role in the out-of-domain scenario (Amazon and FanFic datasets). Similarly, for MultiLUAR trained on Amazon (Figure 3.b), both layers 0 and 6 are more relevant to the in-domain predictions (Amazon instances), while for out-of-domain (Reddit and Fanfiction), also early layers such as 1 to 3 contribute to the predictions. Finally, as for the MultiLUAR model trained on Fanfiction (Figure 3.c), we see a clearer pattern where layer 6 contributes mostly to the in-domain predictions while early layers (0 to 2) predict out-of-domain cases.
6 Discussion and Conclusion
In this paper, we argued that an author’s writing style presents itself at different levels of linguistic granularity. For example, two author texts can be similar (hence likely written by the same author) because they have a similar vocabulary (lexical level) or follow similar structure (discourse level). Therefore, to address the authorship attribution task more effectively, we proposed a new approach that leverages all the representations learned at different layers of the base transformer instead of using only the final layer. This feature of our approach becomes more important when evaluating models out-of-domain since utilizing multiple levels of linguistic granularity enriches the model’s representations, allowing it to perform more robustly.
We evaluated our approach against state-of-the-art baselines on three datasets. Results confirmed how our approach leads to increased performance on the task in the out-of-domain scenario while maintaining relevantly strong performance in the in-domain settings. This gap between our model and state-of-the-art becomes apparent when training is carried out on a more homogeneous dataset like Fanfic and Amazon compared to Reddit.
We then studied what linguistic representations are captured at the different layers of the base transformer (RoBERTa) that our model builds on, and we compared how these representations change in the corresponding layers of our model. Our investigation showed that while RoBERTa’s layers capture more semantics and discourse, the layers of our model capture morphology and syntax relatively better, indicating the importance of these two linguistic properties for the AA task. We further analyze the contribution of each layer in our model towards the final prediction through an ablation study. Across the three trained models, we found a trend of having layers contributing more to the out-of-domain performance than in-domain, such as layers 0 and 1 in MultiLUAR trained on Reddit. Latter layers of all models were found to be essential to the in-domain performance. Overall, our findings demonstrate the importance of leveraging all transformer layers when addressing the authorship attribution task to achieve robust performance in out-of-domain scenarios.
7 Limitations
First, we applied our approach to only one authorship attribution model, which is LUAR Rivera-Soto et al. (2021). Although we tested our hypothesis on three different datasets and showed a consistent behavior, one can potentially demonstrate the robustness of our method by applying it to other transformer-based AA models.
Second, in our analysis of the linguistic competence of each of the layers of our MultiLUAR model, we rely on some probing methods. However, these might not be very reliable tools.
Ethics Statement
We acknowledge that models developed for authorship attribution raise ethical concerns. They can be used to reveal the identity of individuals. While this might be useful for scenarios such as detecting online hate crimes, it could also be used to censor freedom of speech.
Besides, state-of-the-art systems for authorship attribution might suffer from generating false positives, for example, scenarios where a post is wrongly attributed to an individual in a criminal case. Such false positives become more frequent in out-of-domain scenarios. We believe that the research in this paper is a step towards building more robust authorship attribution systems.
Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein
Appendix A Experimental Setup
Model.
For our experiments, we utilize the RoBERTa model (paraphrase-distilroberta) Sanh et al. (2020) of 82.5 M parameters size, including seven additional linear layers, as the backbone for both the MultiLUAR approach and the baseline.
Our implementation follows LUAR’s Rivera-Soto et al. (2021) approach to memory-efficient attention, which replaces the traditional attention mechanism in the transformer architecture. Rabe and Staats (2021) optimized memory usage by chunking queries, keys, and values into smaller segments, thereby reducing memory overhead without sacrificing computational efficiency. Instead of computing the full attention matrix simultaneously, attention is applied in smaller query chunks (size: 32) and key-value chunks (size: 32). The core optimization is achieved through log-space normalization, where the maximum log value is subtracted before exponentiation to prevent overflow. Attention weights and values are then aggregated in chunks, with final normalization performed only at the end. This method reduces the memory footprint while maintaining the standard O(n²) time complexity, achieving a significantly improved O(log n) memory complexity. Additionally, PyTorch’s activation checkpointing is employed, allowing selective recomputation of activations during backpropagation, which further optimizes GPU memory usage.
Training.
Training is conducted using an NVIDIA A100 GPU (40GB memory) to handle the computational demands of large-scale transformer-based models. Our training spans three datasets : Reddit, Fanfiction, and Amazon reviews with a fixed training schedule of 20 epochs per dataset. Specifically, training on the Reddit dataset requires approximately 1 hour per epoch, while Fanfiction and Amazon reviews take 11 minutes and 5 minutes per epoch, respectively.
We do not perform any hyperparameter tuning, opting instead to use the default parameters from the LUAR implementation. Though we experimented with training beyond 20 epochs with some minute performance gains, we adhere to the original LUAR training setup to ensure reproducibility of prior results.
Our training pipeline uses PyTorch Lightning (1.8.5.post0), PyTorch (1.13.1), Transformers (4.14.1), and Scikit-learn (1.2.0), which facilitate efficient model training, evaluation, and data handling.
Datasets Content.
Our study is based on publicly available datasets, which have been widely used in prior research on author attribution and stylistic analysis. These datasets originate from online platforms and may inherently contain offensive or personally identifiable content. However, we do not apply any additional filtering beyond what has been done in previous studies. To ensure ethical usage, we follow the guidelines set by the original sources that published these datasets and acknowledge any potential biases or content-related concerns that may arise.
Additionally, we discuss the licensing terms associated with the datasets used in our experiments. The datasets are provided under open-access licenses, allowing their use for academic research. We ensure compliance with these licenses and properly attribute the sources in our work. Detailed information on dataset licensing is included in the final version of the paper.
Appendix B Ablation Study
B.1 Significance of Layer 0
Our model, MultiLUAR, consists of seven projection layers, where Layer 0 represents the projection of the initial token embeddings from the distillroberta transformer, while Layers 1 to 6 correspond to the projection of its six hidden layers. To demonstrate that, although Layer 0 is not part of the model architecture, it plays a significant role in authorship attribution, we train and evaluate all models on the three datasets without adding a linear projection to Layer 0. The embedding is not incorporated into the loss calculation and similarity metric computation. The results in Table 3 compare the performance of MultiLUAR with and without Layer 0 across different evaluation datasets. We observe that including Layer 0 generally improves performance in out-of-domain evaluation, particularly on the Amazon and Fanfiction datasets, where it enhances R@8 and MRR scores. This suggests that initial token embeddings contain valuable stylistic information that contributes to generalization across domains. However, for in-domain evaluation on Reddit, the model without Layer 0 slightly outperforms the full model as shown in section 5.2.1, indicating that deeper layers are more influential in capturing dataset-specific stylistic patterns. These findings reinforce the importance of leveraging multi-layer representations for robust authorship attribution.
Evaluation Dataset Model Reddit Amazon Fanfic
R@8 MRR R@8 MRR R@8 MRR
Reddit MultiLUAR w [0] 67.87 53.31 26.92 18.03 21.11 14.32
MultiLUAR wø[0] 68.55 53.7 26.76 17.94 19.38 12.73
Amazon MultiLUAR w [0] 72.63 60.74 84.68 72.65 56.69 44.36
MultiLUAR wø[0] 72.14 59.58 84.12 71.78 52.97 40.51
Fanfic MultiLUAR w [0] 43.26 33.96 39.21 29.98 51.78 44.33
MultiLUAR wø[0] 42.84 33.44 38.64 28.65 51.41 42.28
Table 3: Performance comparison of MultiLUAR with and without [0] across different training datasets. Bold values indicate the better performance between the two variations.